### v1

# From Path Coefficients to Targeted Estimands: A Comparison of Structural Equation Models (SEM) and Targeted Maximum Likelihood Estimation (TMLE)

Abstract
Structural Equation Modeling (SEM) has gained popularity in the social sciences and causal inference due to its flexibility in modeling complex relationships between variables and its availability in modern statistical software. To move beyond the parametric assumptions of SEM, this paper reviews targeted maximum likelihood estimation (TMLE), a doubly robust, machine learning-based approach that builds on nonparametric SEM. We demonstrate that both TMLE and SEM can be used to estimate standard causal effects and show that TMLE is robust to model misspecification. We conducted simulation studies under both correct and misspecified model conditions, implementing SEM and TMLE to estimate these causal effects. The simulations confirm that TMLE consistently outperforms SEM under misspecification in terms of bias, mean squared error, and the validity of confidence intervals. We applied both approaches to a real-world dataset to analyze the mediation effects of poverty on access to high school, revealing that the direct effect is no longer significant under TMLE, whereas SEM indicates significance. We conclude with practical guidance on using SEM and TMLE in light of recent developments in targeted learning for causal inference.
keywords:
Structural Equation Modeling, Targeted Learning, Causal Inference, Mediation Analysis, Super Learner
1 Introduction
Structural Equation Modeling [36, SEM;] is a widely used framework in applied statistics, social sciences, psychology, and related fields [52, 62, 48, e.g.,], allowing researchers to model complex interrelationships among observed and latent variables. As a multivariate technique that often incorporates path analysis, SEM also provides an approach to capture mediational mechanisms through direct and indirect effects [39, 46]. In addition to SEM’s adaptability and flexibility, many software implementations are available (e.g., lavaan [59] in R; Stata; and Mplus), which offer further applicability for empirical research.
Numerous research questions in sociology and other fields such as psychology and education could be equivalently answered by estimating the path coefficients within the SEM framework. Under the structural equations and the distributional assumptions on the disturbance terms, parameters are typically estimated by maximizing the likelihood function induced by the model. The subsequent inferences are heavily based on likelihood-based tests or non-parametric approaches such as bootstrap [39, 36]. However, as the true data-generating mechanism typically remains unknown, it is not guaranteed that the specified likelihood encodes it correctly [39]. In the presence of model misspecification, the efficiency and consistency of the estimator and the validity of the corresponding statistical inference are no longer granted, resulting in unreliable estimates and inferences[49, 67].
[…]
Figure 1: A Toy Example of a Path Diagram Depicting Causal Relationships Between Explanatory Variables $X_{1}$ and $X_{2}$ and Response Variable $Y$ with Path Coefficients $\beta_{1}$ and $\beta_{2}$.
Notably, path coefficients and path diagrams (e.g., Figure 1) in SEM can sometimes represent causal effects that are defined under the Neyman–Rubin potential outcomes framework, linking causal inference to parameter estimation [37, 55, e.g.,]. However, the equivalence between path coefficients and causal effects essentially depends on the correct model specification. To address the limitations of linearity and reduce the risk of model misspecification, nonparametric structural equation models [56, NPSEM;] use nonparametric equations to represent relationships between variables, reformulating path diagrams as Directed Acyclic Graphs (DAGs; Figure 2). Using NPSEM and DAG, causal effects are identified as functionals of the joint distribution rather than path coefficients. Built on NPSEM, Targeted Maximum Likelihood (Loss) Estimation (TMLE), introduced by van der Laan and Rubin [64], provides a modern, semiparametric, machine learning-based framework for causal inference that is doubly robust and efficient under certain conditions.
$X_{1}$$X_{2}$$Y$
Figure 2: A Toy Example of a DAG Depicting Causal Relationships Between Variables $X_{1}$ and $X_{2}$ and outcome $Y$.
Upon identifying a causal effect from NPSEM, TMLE then involves two steps: an initial data-adaptive estimation followed by an adjustment that targets the estimation equation of the efficient influence function (EIF) [44]. TMLE is often coupled with the Super Learner algorithm [65]—an ensemble method that combines multiple learners to enhance predictions. Under regularity conditions, TMLE achieves asymptotic efficiency and hence provides a valid uncertainty measurement through the influence function [44, 64], offering computational advantages over bootstrap methods while preserving interpretability.
TMLE offers a robust alternative to traditional parametric SEM, particularly in mitigating the risks of model misspecification across diverse causal inference settings and for estimating various causal effects of interest. We provide a concise review of the use of SEM and TMLE in commonly considered causal inference problems and demonstrate that they share a common conceptual foundation. We aim to show sociologists and applied researchers that TMLE remains reliable through its double robustness and incorporation of the super learner algorithm, and outperforms traditional linear SEM in various settings.
The structure of the paper is as follows: In Section 2, we provide a concise review of the overall frameworks of SEM and TMLE. In Section 3, we define fundamental causal effects and the assumptions required for their identification. We then illustrate the potential applications of SEM and TMLE in estimating these causal effects and conducting causal mediation analysis. In Section 4, we conduct simulation studies under correctly specified models and various violated assumptions, including omitted interaction terms, non-linear relationships, and non-normality, to evaluate and compare the performance of TMLE and SEM. In Section 5, we present real data analyses using both SEM and TMLE, studying the mediational effects of poverty on access to high school. We end this paper by giving some concluding remarks in Section 6.
2 Review of SEM and TMLE
2.1 Structural Equation Modeling
[39] offers a concise review of SEM, from which the following overview is primarily adapted, with minor modifications for notation and clarity. The standard SEM consists of two parts: the measurement submodel and the structural submodel. In general, the structural equation is defined as
$\boldsymbol{\eta}=\boldsymbol{\mu}+\boldsymbol{\beta}\boldsymbol{\eta}+\boldsymbol{\zeta},$
where $\boldsymbol{\mu}\in\mathbb{R}^{k}$ is the vector of latent intercepts and $\boldsymbol{\eta}\in\mathbb{R}^{k}$ is the vector of latent factor scores, $\boldsymbol{\beta}$ is a $k\times k$ matrix representing the regression parameters among the latent factors, and $\boldsymbol{\zeta}$ is a $k\times 1$ vector of normally distributed disturbances with mean vector $\boldsymbol{0}$ and covariance matrix $\boldsymbol{\Psi}$. The measurement model links the latent variables to the observed outcomes
$\boldsymbol{y}=\boldsymbol{\nu}+\boldsymbol{\Lambda}\boldsymbol{\eta}+\boldsymbol{\epsilon},$
where $\boldsymbol{y}\in\mathbb{R}^{p}$ is the vector of observed variables, $\boldsymbol{\nu}$ is a $p\times 1$ vector of measurement intercepts, $\boldsymbol{\Lambda}$ is a $p\times k$ matrix of factor loadings relating $\boldsymbol{y}$ to $\boldsymbol{\eta}$, and $\boldsymbol{\epsilon}$ is a $p\times 1$ vector of normally distributed measurement errors with mean $\boldsymbol{0}$ and covariance matrix $\boldsymbol{\Sigma}_{\epsilon}$. We also assume that $\mathbb{E}[\boldsymbol{\zeta}]=\mathbb{E}[\boldsymbol{\epsilon}]=\boldsymbol{0}$ and $\mathrm{Cov}(\boldsymbol{\zeta},\boldsymbol{\epsilon})=\boldsymbol{0}$. Substituting the structural equation into the measurement model, we may alternatively express the model as
$\boldsymbol{y}=\boldsymbol{\nu}+\boldsymbol{\Lambda}\boldsymbol{B}\boldsymbol{\mu}+\boldsymbol{\Lambda}\boldsymbol{B}\boldsymbol{\zeta}+\boldsymbol{\epsilon},$
where $\boldsymbol{B}=(I_{k}-\boldsymbol{\beta})^{-1}$. This formulation implies that the mean and covariance of $\boldsymbol{y}$ are
$\mathbb{E}[\boldsymbol{y}]=\boldsymbol{\nu}+\boldsymbol{\Lambda}\boldsymbol{B}\boldsymbol{\mu}:=\boldsymbol{\mu}_{y},\quad\mathrm{Cov}(\boldsymbol{y})=\boldsymbol{\Lambda}\boldsymbol{B}\boldsymbol{\Psi}\boldsymbol{B}^{\top}\boldsymbol{\Lambda}^{\top}+\boldsymbol{\Sigma}_{\epsilon}:=\boldsymbol{\Sigma}_{y}.$
As mentioned, the parameters are generally estimated by maximizing the induced likelihood. Given the normality of the disturbances, the joint log-likelihood for a sample of size $n$ (with sample mean $\bar{\boldsymbol{y}}$ and sample covariance matrix $\boldsymbol{S}$) can be written as
$\ell(\boldsymbol{\theta})=-\frac{n}{2}\left(\log|\boldsymbol{\Sigma}_{y}|+\text{tr}(\boldsymbol{\Sigma}_{y}^{-1}\boldsymbol{S})+(\bar{\boldsymbol{y}}-\boldsymbol{\mu}_{y})^{\top}\boldsymbol{\Sigma}_{y}^{-1}(\bar{\boldsymbol{y}}-\boldsymbol{\mu}_{y})+p\log(2\pi)\right),$
where $\boldsymbol{\theta}$ is the collection of all free parameters. Neglecting constants, we may estimate $\boldsymbol{\theta}$ by minimizing the negative log-likelihood function
$\hat{\boldsymbol{\theta}}=\arg\min_{\boldsymbol{\theta}}-\ell(\boldsymbol{\theta})=\arg\min_{\boldsymbol{\theta}}\log|\boldsymbol{\Sigma}_{y}|+\text{tr}(\boldsymbol{\Sigma}_{y}^{-1}\boldsymbol{S})+(\bar{\boldsymbol{y}}-\boldsymbol{\mu}_{y})^{\top}\boldsymbol{\Sigma}_{y}^{-1}(\bar{\boldsymbol{y}}-\boldsymbol{\mu}_{y}).$
Under regularity conditions, the properties of MLE ensure the asymptotic distribution of $\sqrt{n}(\hat{\boldsymbol{\theta}}_{\text{ML}}-\boldsymbol{\theta})$ is multivariate normal with mean $\boldsymbol{0}$ and covariance matrix given by the inverse of the Fisher information matrix, enabling statistical inference such as Wald tests for parameter significance and likelihood ratio tests for model fit. It is possible that the desired estimation involves some nonlinear transformations of the path coefficients. In such a case, the uncertainty is usually captured by the Delta method or non-parametric techniques like bootstrap [37].
2.2 Targeted Learning and the SuperLearner
2.2.1 An Overview of Targeted Learning
In contrast to parametric SEM, Targeted Maximum Likelihood(Loss) Estimation is a modern, semi-parametric estimation framework developed to provide an efficient and robust estimation of causal effects and other statistical parameters in complex data settings. An NPSEM is generally specified first for describing the causal relations and identifying the causal parameters in terms of the joint distribution. Then, the TMLE involves two main steps: an initial estimation of outcome functions using any suitable method, and a targeting step that updates these estimates to improve the estimation of the parameter of interest. This targeting is done through a fluctuation submodel and solving the estimation equation of the efficient influence function (EIF), which plays a key role in ensuring local efficiency. The resulting TMLE estimator is asymptotically linear and normally distributed under mild regularity conditions, allowing for valid inference via standard error estimation and confidence intervals. Following [45], we assume that a semi-parametric statistical model has been identified $\mathcal{M}$ with a true but unknown distribution $F_{0}\in\mathcal{M}$. The interested parameter is defined as $\Psi(F_{0}),\Psi(\cdot):\mathcal{M}\rightarrow\mathbb{R}$, and with $O_{1},O_{2},\cdots,O_{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}F_{0}$ observations from $F_{0}$. We further need that $\Psi(F_{0})=\Psi(Q_{0})$ depends on only $Q_{0}=Q(F_{0})$, a part of $F_{0}$, and $g_{0}$ is a nuisance parameter from some orthogonal factorization.
The first step in TMLE is to get an initial estimate of $Q_{0}$, denoting $Q_{n}^{0}$. This would require a specification of the loss function $\mathcal{L}_{Q}$ such that
$Q_{0}=\arg\min_{Q\in\mathcal{Q}}\mathbb{E}_{O\sim F_{0}}\mathcal{L}_{Q}(O),\quad\mathcal{Q}:=\{Q(F):F\in\mathcal{M}\},$
where common choices are negative log-likelihood and squared loss depending on the outcome. The estimation can then be obtained by solving
$Q^{0}_{n}=\arg\min_{Q\in\mathcal{Q}}\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{Q}(O_{i}),$
which would yield $Q^{0}_{n}(O_{i}),\forall i\in\{1,2,\cdots,n\}$ by plugging-in the observations. The second step in TMLE is the targeting step, which first requires an estimator $g_{n}$ of the nuisance parameter $g$. This can also be done via parametric regression or loss-based learning, depending on the true functional form. Suppose we have obtained $g_{n}$, one needs to propose a parametric fluctuation $Q_{n,g}^{1}(\epsilon)$ that satisfies
$\dfrac{\mathrm{d}}{\mathrm{d}\epsilon}\mathcal{L}\left(Q_{n,g}^{1}(\epsilon)\right)(O)\big|_{\epsilon=0}=D^{*}(Q_{n}^{0},g)(O),$
where the $D^{*}$ on the RHS is the EIF of $\Psi:\mathcal{M}\rightarrow\mathbb{R}$ at $F_{0}$. By solving the fluctuation parameter based on the observations $O_{i}$
$\epsilon_{n}^{1}=\arg\min_{\epsilon}\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}\left(Q_{n,g_{n}}^{1}(\epsilon)\right)(O_{i}),$
we can update the initial estimate as $Q_{n,g}^{1}(\epsilon_{n}^{1})$. This might be iterated until convergence. Specifically, in the $i^{th}$ iteration, we do
$\epsilon_{n}^{i}=\arg\min_{\epsilon}\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}\left(Q_{n,g_{n}}^{i}(\epsilon)\right)(O_{i}),\quad\text{where }\dfrac{\mathrm{d}}{\mathrm{d}\epsilon}\mathcal{L}\left(Q_{n,g}^{i}(\epsilon)\right)(O)\big|_{\epsilon=0}=D^{*}(Q_{n}^{i-1},g)(O).$
The updates from the last iteration, denoting $Q_{n}^{*}$, should satisfy the estimation equation of the empirical efficient influence function
$\dfrac{1}{n}\sum_{i=1}^{n}D^{*}(Q_{n}^{*},g_{n})(O_{i})=0,$
which would result in an estimator $\Psi(Q_{n}^{*})$ that attains the asymptotic efficiency, i.e.
$\sqrt{n}\left(\Psi(Q_{n}^{*})-\Psi(F_{0})\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}D^{*}(O_{i})+o_{p}(1).$
Thus, by the linearity, the limiting distribution of $\Psi(Q^{*}_{n})$, as $n\rightarrow\infty$, is
$\sqrt{n}\left(\Psi(Q_{n}^{*})-\Psi(F_{0})\right)\rightarrow^{d}\mathcal{N}(0,\mathrm{Var}(D^{*})),$
which also yields a valid confidence set. This framework has been adapted to many estimands that admit the asymptotic linear estimation (e.g. [45, 47, 61]).
2.2.2 The Super Learner
As mentioned, the efficiency and consistency of TMLE depend heavily on the initial estimates of $Q_{n}^{0}$ and $g_{n}$. Though machine learning algorithms could be applied instead of the parametric regression, we by no means should rely on a single algorithm, as none is optimal for all datasets. It is worth mentioning that the estimation of $Q_{n}^{0}$ and $g_{n}$ can be done non-parametrically via the Super Learner [63], an ensemble learning technique that integrates modern machine learning algorithms. Consider a dataset $\{(X_{i},Y_{i})\}_{i=1}^{n}$, where $X_{i}\in\mathbb{R}^{p}$ are features and $Y_{i}\in\mathbb{R}$ are outcomes, jointly drawn from an unknown distribution $F_{0}\in\mathcal{M}$. Recall that the initial estimates in TMLE are equivalent to constructing a function $f(X)$ that minimizes the expected risk
$R(f)=\mathbb{E}_{F_{0}}[\mathcal{L}(Y,f(X))],$
where $\mathcal{L}$ is a loss function, such as squared error $(Y-f(X))^{2}$ for regression or log-loss for classification. Instead of the simple ensemble, Super Learner constructs a predictor $f_{\text{SL}}(X)=\sum_{m=1}^{M}\hat{w}_{m}f_{m}(X)$, where $\{f_{m}\}_{m=1}^{M}$ is the predictions from user-specified base learners, and weights $\hat{w}_{m}\geq 0$, $\sum_{m=1}^{M}\hat{w}_{m}=1$, are determined by minimizing the risk. In particular, to estimate weights, Super Learner employs $V$-fold cross-validation. The dataset is partitioned into V mutually exclusive subsets $\{D_{v}\}_{v=1}^{V}$. For each fold $v$, each base learner is trained on the training set $D_{-v}=D\setminus D_{v}$, producing predictions on validation set $Z_{i,m}=f_{m}^{-v}(X_{i})$ for $i\in D_{v}$. These predictions form a level-one data matrix $Z\in\mathbb{R}^{n\times M}$. A meta-learner, typically non-negative least squares (NNLS), solves
$\hat{w}=\arg\min_{w\in\mathcal{W}}\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}\left(Y_{i},\sum_{m=1}^{M}w_{m}Z_{i,m}\right),$
where $\mathcal{W}=\{w:w_{m}\geq 0,\sum_{m=1}^{M}w_{m}=1\}$. The resulting super predictor admits the Oracle property under mild conditions, ensuring it asymptotically achieves the optimal weighted combination of base learners.
Figure 3: Illustration of the Super Learner algorithm flow. Adapted from [63].
Numerous modern machine learning algorithms are included in the current SuperLearner implementation in R, including Random Forest, Support Vector Machine, eXtreme Gradient Boosting, and Artificial Neural Network, which enhances its versatility and applicability. Notably, the choices of base learners and the meta learner play a key role in the final estimation, and have to be considered based on the specific tasks involved [57].
3 SEM and TMLE approach to causal inference
When considering SEM for causal inference, the commonly considered parameters are the Average Treatment Effect (ATE), Conditional Average Treatment Effect (CATE), and mediational effects [46, 37]. In this section, we consider these causal effects and show that both SEM and TMLE are applicable for estimation.
3.1 On Average Treatment Effects
Suppose that we observed independent and identical copies of $\{(Y_{i},A_{i},\boldsymbol{X}_{i})\}_{i=1}^{n}$, where $Y_{i}$ is the outcome measurement, $A_{i}$ is the binary treatment, and $\boldsymbol{X}_{i}=(X_{i1},X_{i2},\cdots X_{ip})$ is the pre-treatment covariates take value in $\mathcal{X}\subseteq\mathbb{R}^{p}$. We let $Y(1)$ denote the outcome when treatment is assigned, i.e. $A=1$, and $Y(0)$ denotes the outcome when $A=0$. ATE is then defined to be the expected difference between potential outcomes $Y(1)$ and $Y(0)$, i.e.
$\psi_{\textup{ATE}}=\mathbb{E}[Y(1)-Y(0)].$
To capture the heterogeneity, the Conditional Average Treatment Effect (CATE) for some sub-group $B\subset\mathcal{X}$ is defined similarly
$\psi_{\textup{CATE}}=\mathbb{E}[Y(1)-Y(0)|\boldsymbol{X}\in B],\quad B\subset\mathcal{X}.$
Three essential causal assumptions for efficient identification and estimation of the causal effect from accessible data are (i) Stable Unit Treatment Value Assumption (SUTVA) or the consistency assumption: $Y=A\cdot Y(1)+(1-A)\cdot Y(0)$; (ii) no unmeasured confounders: $\mathbb{P}(A=1|Y(1),Y(0),\boldsymbol{X})=\mathbb{P}(A=1|\boldsymbol{X})$; and (iii) positivity: $\exists\epsilon>0\text{such that }\mathbb{P}(A=1|\boldsymbol{X}=\boldsymbol{x})>\epsilon$ for almost all $\boldsymbol{x}\in\mathcal{X}$, without which the ATE might not be identifiable [58].
3.1.1 The SEM approach to ATE
Suppose that the data-generating mechanism is encoded by the structural linear causal models [55]
$(U_{\boldsymbol{X}},U_{A},U_{Y})\sim F_{U}$
$\boldsymbol{X}=U_{\boldsymbol{X}}$
$A=g_{A}(\boldsymbol{X},U_{A})$
$Y=\gamma A+\boldsymbol{\beta}^{\top}\boldsymbol{X}+U_{Y},$
where $(U_{\mathbf{X}},U_{A},U_{Y})$ are unobserved but mutually uncorrelated exogenous variables and $g_{A}$ is known (e.g., logistic). For concreteness, we assume linearity, but extensions to other parametric forms can be easily adapted. Under the causal assumptions that we specified above, estimating $\psi_{\text{ATE}}=\mathbb{E}(Y(1))-\mathbb{E}(Y(0))$ is equivalent to estimating the path coefficient $\gamma$:
$\psi_{\text{ATE}}$ $=\mathbb{E}_{\boldsymbol{X}}\left[\mathbb{E}[Y(1)|\boldsymbol{X}]-\mathbb{E}[Y(0)|\boldsymbol{X}]\right]$
$=\mathbb{E}_{\boldsymbol{X}}\left[\mathbb{E}[Y|A=1,\boldsymbol{X}]-\mathbb{E}[Y|A=0,\boldsymbol{X}]\right]$
$=\int_{\mathcal{X}}\mathbb{E}[Y|A=1,\boldsymbol{X}]-\mathbb{E}[Y|A=0,\boldsymbol{X}]\mathrm{d}F_{\boldsymbol{X}}(\boldsymbol{x})$
$=\int_{\mathcal{X}}\left(\gamma+\boldsymbol{x}\boldsymbol{\beta}-\boldsymbol{x}\boldsymbol{\beta}\right)\mathrm{d}F_{\boldsymbol{X}}(\boldsymbol{x})=\int_{\mathcal{X}}\gamma\mathrm{d}F_{\boldsymbol{X}}(\boldsymbol{x})=\gamma,$
[…]
$(\hat{\gamma},\hat{\boldsymbol{\beta}})=\arg\max_{\gamma,\beta}\ell(O;\gamma,\boldsymbol{\beta})=\arg\max_{\gamma,\beta}\frac{1}{n}\sum_{i=1}^{n}\log\phi(\mu_{i}(A_{i},\boldsymbol{X}_{i}),\sigma^{2})$
where $\ell$ denotes the log-likelihood function with $\phi(\mu,\tau^{2})$ denotes the normal density with mean $\mu$ and variance $\tau^{2}$, and $O$ represents the observed data, or as part of a generalized SEM with the logit link. It is unbiased, consistent, efficient, and asymptotically normal under correct model specification [66]. The statistical inference regarding the ATE $H_{0}:\gamma=0$, in a SEM context, is usually done via the Wald test and the likelihood ratio test.
A more complex scenario arises when the model incorporates interaction terms between the treatment $A$ and certain covariates, reflecting heterogeneity in treatment effects. For concreteness, consider the case where the treatment interacts with the covariate $X_{ik}$:
$Y_{i}=\gamma A_{i}+\boldsymbol{\beta}^{\top}\boldsymbol{X}_{i}+\tau A_{i}X_{ik}+U_{Y}.$
In this setting, the causal effect varies across subgroups defined by different values of $X_{ik}$. We thus consider the conditional average treatment effect (CATE) for a subgroup where $\boldsymbol{X}\in B\subset\mathcal{X}$:
$\psi_{\text{CATE}}$ $=\mathbb{E}[Y(1)\mid\boldsymbol{X}\in B]-\mathbb{E}[Y(0)\mid\boldsymbol{X}\in B]$
$=\mathbb{E}[Y\mid A=1,\boldsymbol{X}\in B]-\mathbb{E}[Y\mid A=0,\boldsymbol{X}\in B]$
$=\int_{B}\mathbb{E}[Y\mid A=1,\boldsymbol{x}]-\mathbb{E}[Y\mid A=0,\boldsymbol{x}]\,\mathrm{d}F_{\boldsymbol{X}|\boldsymbol{X}\in B}(\boldsymbol{x})$
$=\int_{B}(\gamma+\boldsymbol{\beta}^{\top}\boldsymbol{x}+\tau x_{k}-\boldsymbol{\beta}^{\top}\boldsymbol{x})\,\mathrm{d}F_{X}(\boldsymbol{x}\mid\boldsymbol{X}\in B)=\gamma+\tau\mathbb{E}[X_{k}\mid\boldsymbol{X}\in B].$
If $X_{k}=x_{k}$ is constant in $B$, the CATE simplifies to $\gamma+\tau x_{k}$. Estimation follows similarly, with asymptotic normality preserved under linear transformations, and the inference remains valid up to a minor modification [66].
3.1.2 The targeted learning approach to ATE
However, the identification of causal effects as path coefficients and the consistency of the estimators rely on correct functional form specification, and misspecification can introduce bias[39, 67]. In contrast to parametric SEM, considering the following NPSEM [55]
$(U_{\boldsymbol{X}},U_{A},U_{Y})\sim F_{U}$
$\boldsymbol{X}=f_{\boldsymbol{X}}(U_{\boldsymbol{X}})$
$A=f_{A}(\boldsymbol{X},U_{A})$
$Y=f_{Y}(A,\boldsymbol{X},U_{Y}),$
where functions $f_{A},f_{\boldsymbol{X}},f_{Y}$ are no longer restricted to be linear or known. Hence, the wanted estimate is no longer the path coefficients but a mapping from the statistical model to the real line. TMLE for ATE leverages the orthogonal factorization of the likelihood function induced by the NPSEM [44]
$\mathcal{L}(Y,A,\boldsymbol{X})=\mathbb{P}(Y|A,\boldsymbol{X})\mathbb{P}(A|\boldsymbol{X})\mathbb{P}(\boldsymbol{X}).$
We define $Q(A,\boldsymbol{X})=\mathbb{E}(Y|A,\boldsymbol{X})$ and $g(A|\boldsymbol{X})=\mathbb{P}(A|\boldsymbol{X})$, where $Q(A,\boldsymbol{X})$ can be estimated from the observed data and the $g(A|\boldsymbol{X})$ is nuisance parameter that will be used in the subsequent targeting step [44]. For concreteness, we assume the measurement $Y$ is continuous, while other types, such as binary, can also be accommodated with minor modifications [53]. For an unbounded continuous outcome, it is generally recommended to scale $Y$ with the min-max transformation [42, 64]
$\tilde{Y}_{j}=\dfrac{Y_{j}-\min_{i}Y_{i}}{\max_{i}Y_{i}-\min_{i}Y_{i}}\in[0,1],$
followed by the inverse transformation after the point and set estimations are completed. The first step in implementing TMLE is to obtain an initial estimate of $\mathbb{E}(Y|A,\boldsymbol{X})$. Since the outcome is bounded in $[0,1]$, one could use logistic regression or machine learning algorithms for estimation, but incorporating cross-validation to avoid overfitting is crucial and generally recommended [45]. Predictions from the fitted or trained model yield $Q_{n}^{0}(A=1,\boldsymbol{X}_{i})$ and $Q_{n}^{0}(A=0,\boldsymbol{X}_{i}),\forall i\in\{1,2,\cdots,n\}$. The targeting step begins with estimating the propensity score, $g(A\mid\boldsymbol{X})$. The clever covariate $H(g,A,\boldsymbol{X})$, derived from the efficient influence function, is defined to be
$H(g_{n},A,\boldsymbol{X})=\frac{A}{g_{n}(A=1|\boldsymbol{X})}-\dfrac{1-A}{g_{n}(A=0|\boldsymbol{X})},$
where $g_{n}$ denotes the estimated propensity score. To complete the targeting step, we consider a fluctuation functional $Q_{n,g}^{*}(\epsilon)=\text{expit}(\text{logit}(Q_{n}^{0}(A,\boldsymbol{X}_{i}))+\epsilon H(g,A,\boldsymbol{X}_{i})),$ where expit($z$) = $\frac{1}{1+e^{-z}}$ is the inverse logit. The targeting step can be seen as solving the estimation equation of the efficient influence function
$\frac{1}{n}\sum_{i=1}^{n}D^{*}(O_{i},Q_{n}^{*}(\epsilon,A_{i},\boldsymbol{X}_{i}),g_{n})=0$
where the EIF for ATE is
$D^{*}(O_{i},Q_{n}^{*}(\epsilon,A_{i},\boldsymbol{X}_{i}),g_{n})=H(g_{n},A,\boldsymbol{X}_{i})(Y_{i}-Q_{n}^{*}(\epsilon,A_{i},\boldsymbol{X}_{i}))+Q_{n}^{*}(\epsilon,1,\boldsymbol{X}_{i})-Q_{n}^{*}(\epsilon,0,\boldsymbol{X}_{i})-\psi.$
The fluctuation parameter $\epsilon$ can be equivalently estimated using the logistic regression
$\text{logit}(Q_{n}(A,\boldsymbol{X}))=\text{logit}(Q_{n}^{0}(A,\boldsymbol{X}))+\epsilon H(A,\boldsymbol{X}),$
and where $\text{logit}(Q_{n}^{0}(A,\boldsymbol{X}))$ serves as an offset term and $\hat{\epsilon}$ is estimated via maximizing the likelihood. Upon estimating the fluctuation parameter, we may update the initial estimate to be
$Q_{n}^{*}(A,\boldsymbol{X}_{i})=\text{expit}(\text{logit}(Q_{n}^{0}(A,\boldsymbol{X}_{i}))+\hat{\epsilon}H(g_{n},A,\boldsymbol{X}_{i})).$
Subsequently, the final TMLE estimation can be obtained by the substitution estimation
$\hat{\psi}_{\text{ATE}}^{\text{TMLE}}=\frac{1}{n}\left(\sum_{i=1}^{n}Q_{n}^{*}(A=1,\boldsymbol{X}_{i})-Q_{n}^{*}(A=0,\boldsymbol{X}_{i})\right).$
It has been established that if either $Q^{0}_{n}(A,\boldsymbol{X})$ or $g_{n}(A\mid\boldsymbol{X})$ is consistent, the resulting $\hat{\psi}^{\operatorname{TMLE}}_{\operatorname{ATE}}$ is a consistent estimator for $\psi_{\operatorname{ATE}}$ [64]. To assess the uncertainty of the TMLE, its variance can be estimated using the efficient influence function:
$\hat{D}(O_{i})=H(g_{n},A_{i},\mathbf{X}_{i})(Y_{i}-Q_{n}^{*}(A_{i},\mathbf{X}_{i}))+Q_{n}^{*}(1,\mathbf{X}_{i})-Q_{n}^{*}(0,\mathbf{X}_{i})-\hat{\psi}_{\text{ATE}}^{\text{TMLE}},$
where $\widehat{\mathrm{Var}}(\hat{\psi}_{\text{ATE}}^{\text{TMLE}})=\frac{1}{n}\sum_{i=1}^{n}[\hat{D}(O_{i})]^{2}/n$. Under regularity conditions, as $n\to\infty$,
$\sqrt{n}(\hat{\psi}^{\operatorname{TMLE}}_{\operatorname{ATE}}-\psi_{\operatorname{ATE}})\to^{d}\mathcal{N}(0,\operatorname{Var}(D^{*}))$
allowing a 95% confidence interval to be constructed as $\hat{\psi}^{\operatorname{TMLE}}_{\operatorname{ATE}}\pm z_{0.975}\hat{\sigma}/\sqrt{n}$ [44]. It has been shown that if both the initial estimate $Q_{n}^{0}$ and the estimated propensity score $g_{n}$ are consistent, then it attains the efficiency [64].
Many tutorials are available for implementing the TMLE for ATE with either a continuous or a binary outcome [42, 53, e.g.,], while the R package tmle also provides a built-in function that can integrate with the SuperLearner [43]. The steps for estimating CATE via TMLE can be done in strata, where the package tmle3 provides a built-in function to achieve this [38].
3.2 Causal Mediation Analysis
Causal mediation analysis is a statistical framework used to understand how a treatment $A$ affects an outcome $Y$ by decomposing the total causal effect into a direct effect (the part not operating through an intermediate variable) and an indirect effect (the part that operates through a mediator, $M$). This helps disentangle mechanisms in causal pathways, common in fields like epidemiology, psychology, and social sciences. Suppose we have continuous outcome $Y$, a binary treatment or exposure $A\in\mathcal{A}=\{0,1\}$, baseline covariates $\boldsymbol{W}=(W_{1},W_{2},\cdots,W_{p})$ take values in $\mathcal{W}\subseteq\mathbb{R}^{p}$ and a continuous mediator $M$ with possible values $\mathcal{M}\subseteq\mathbb{R}$. We use $Y(a,m)$ to denote the outcome when the treatment $A=a$ and the mediator $M=m$; $M(a)$ represents the mediator with $A=a$. We would make the composition assumption, which states that $Y(a,M(a))=Y(a)$ for $a\in\{0,1\}$ [41]. The ATE in the context of mediation analysis can then be defined similarly
$\psi_{\text{ATE}}=\mathbb{E}(Y(1,M(1))-Y(0,M(0)))=\mathbb{E}(Y(1)-Y(0)),$
but the direct and indirect effects require extra specification. The Natural Direct Effect(NDE) is the effect of $A$ on $Y$ if the mediator $M$ were fixed to its natural value under the control condition $A=0$. Statistically, it can be defined as
$\text{NDE}=\mathbb{E}[Y(1,M(0))-Y(0,M(0))],$
while the Natural Indirect Effect(NIE), as a counterpart, is defined to be
$\text{NIE}=\mathbb{E}[Y(1,M(1))-Y(1,M(0))].$
Under the composition assumption, we could decomposite $\psi_{\text{ATE}}=\text{NIE}+\text{NDE}$. To identify the NIE and NDE, we need several assumptions in addition to the standard consistency assumption: (i) sequential ignorability: $A\perp Y(a,m)|\boldsymbol{W}$ and $M\perp Y(a,m)|\boldsymbol{W},A$ for all $a\in\mathcal{A},m\in\mathcal{M}$; (ii) no treatment-mediator confounding: $A\perp M(a)|\boldsymbol{W}$ for all $a\in\mathcal{A}$; and (iii) cross-world independence: $Y(a,m)\perp M(a^{\prime})|\boldsymbol{W}$ for all $a,a^{\prime}\in\mathcal{A},m\in\mathcal{M}$ [41]. The positivity assumption here is extended to be: $\exists\epsilon_{a}>0\text{ such that }\mathbb{P}(A=a\mid\boldsymbol{W}=\boldsymbol{w})>\epsilon_{a}$ $\forall a\in\mathcal{A},\boldsymbol{w}\in\mathcal{W}$ and $\exists\epsilon_{m}>0\text{ such that }\mathbb{P}(M=m\mid A=a,\boldsymbol{W}=\boldsymbol{w})>\epsilon_{m}$ $\forall m\in\mathcal{M},a\in\mathcal{A},\boldsymbol{w}\in\mathcal{W}$. When these assumptions are satisfied, the NDE and NIE can be non-parametrically identified [41, 40]. In particular,
$\text{NDE}=\mathbb{E}_{\boldsymbol{W}}\left[\mathbb{E}_{M|A=0,\boldsymbol{W}}\left(\mathbb{E}(Y|A=1,M,\boldsymbol{W})-\mathbb{E}(Y|A=0,M,\boldsymbol{W})\right)\right]$
$\text{NIE}=\mathbb{E}_{\boldsymbol{W}}\left[\mathbb{E}_{M|A=1,\boldsymbol{W}}(\mathbb{E}(Y|A=1,M,\boldsymbol{W}))-\mathbb{E}_{M|A=0,\boldsymbol{W}}(\mathbb{E}(Y|A=1,M,\boldsymbol{W}))\right].$
In the following subsections, we will assume that these assumptions hold and show the analytic equivalence between SEM and TMLE when the model is correctly specified.
3.2.1 Estimating NDE and NIE via SEM
A commonly used approach for simply the conditional expectation in identifying NIE and NDE is encoding it as a parametric SEM[40]. Suppose our continuous mediator and continuous outcomes are specified under the following structural equations
$U=(U_{\boldsymbol{W}},U_{A},U_{M},U_{Y})\sim P_{U}$
$\boldsymbol{W}=U_{\boldsymbol{W}};$
$A=g_{A}(\boldsymbol{W},U_{A})$
$M=\alpha A+\boldsymbol{\Gamma}^{\top}\boldsymbol{W}+U_{M};$
$Y=\gamma A+\beta M+\boldsymbol{\Theta}^{\top}\boldsymbol{W}+U_{Y},$
where $U$ is the unobserved and uncorrelated exogenous variables, and $g_{A}$ is some known function. We further assume that $M$ has a probability density function $f_{M}$ with respect to some dominating measure and discrete $\boldsymbol{W}$ takes values in some state space $\mathcal{W}\subset\mathbb{R}^{d}$. Since the parameteric assumptions are specified here, the direct and indirect effects can be derived in a closed form [46, 41].
$\boldsymbol{W}$AAMMYY
Figure 4: Causal directed acyclic graph (DAG) depicting relationships among baseline covariates $\boldsymbol{W}$, exposure $A$, mediator $M$, and outcome $Y$.
In particular, we may identify the NDE to be
NDE $=\sum_{\boldsymbol{w}\in\mathcal{W}}\int_{m\in\mathcal{M}}\left[\mathbb{E}(Y|A=1,M=m,\boldsymbol{W}=\boldsymbol{w})-\mathbb{E}(Y|A=0,M=m,\boldsymbol{W}=\boldsymbol{w})\right]$
$\times f(M=m|A=0,\boldsymbol{W}=\boldsymbol{w})\mathrm{d}m\mathbb{P}(\boldsymbol{W}=\boldsymbol{w})$
$=\sum_{\boldsymbol{w}\in\mathcal{W}}\int_{m\in\mathcal{M}}\left(\gamma+\beta m+\boldsymbol{\Theta}^{\top}\boldsymbol{w}-\beta m-\boldsymbol{\Theta}^{\top}\boldsymbol{w}\right)\times f(M=m|A=0,\boldsymbol{W}=\boldsymbol{w})\mathrm{d}m\mathbb{P}(\boldsymbol{W}=\boldsymbol{w})$
$=\sum_{\boldsymbol{w}\in\mathcal{W}}\int_{m\in\mathcal{M}}\gamma\times f(M=m|A=0,\boldsymbol{W}=\boldsymbol{w})\mathrm{d}m\mathbb{P}(\boldsymbol{W}=\boldsymbol{w})=\sum_{\boldsymbol{w}\in\mathcal{W}}\gamma\mathbb{P}(\boldsymbol{W}=\boldsymbol{w})=\gamma$
The NIE can be identified similarly by
NIE $=\sum_{\boldsymbol{w}\in\mathcal{W}}\int_{m\in\mathcal{M}}\mathbb{E}(Y|A=1,M=m,\boldsymbol{W}=\boldsymbol{w})\times\{f(M=m|A=1,\boldsymbol{W}=\boldsymbol{w})$
$-f(M=m|A=0,\boldsymbol{W}=\boldsymbol{w})\}\mathrm{d}m\mathbb{P}(\boldsymbol{W}=\boldsymbol{w})$
$=\sum_{\boldsymbol{w}\in\mathcal{W}}\int_{m\in\mathcal{M}}(\gamma+\beta m+\boldsymbol{\Theta}^{\top}\boldsymbol{w})\times\{f(M=m|A=1,\boldsymbol{W}=\boldsymbol{w})$
$-f(M=m|A=0,\boldsymbol{W}=\boldsymbol{w})\}\mathrm{d}m\mathbb{P}(\boldsymbol{W}=\boldsymbol{w})$
$=\sum_{\boldsymbol{w}\in\mathcal{W}}\beta(\mathbb{E}(M|A=1,\boldsymbol{W}=\boldsymbol{w})-\mathbb{E}(M|A=0,\boldsymbol{W}=\boldsymbol{w}))\mathbb{P}(\boldsymbol{W}=\boldsymbol{w})$
$=\sum_{\boldsymbol{w}\in\mathcal{W}}\beta\alpha\mathbb{P}(\boldsymbol{W}=\boldsymbol{w})=\beta\alpha.$
This shows that under correct model specification, estimating the NIE and NDE is equivalent to estimating the path coefficients in SEM. We may also identify it using the path analysis and coefficients; there are two paths from $A$ to $Y$: $A\rightarrow M\rightarrow Y$ and $A\rightarrow Y$, where the former passes $M$, which would be the indirect effect and the latter would be the direct effect. Given the independence of $U_{M}$ and $U_{Y}$, $Y$ would follow a normal distribution with the model-induced distribution parameters. Hence, the parameter estimation could be done by maximizing the model-induced likelihood with the following Wald test for inference. For NIE, the inference is commonly done via the Delta method or via Bootstrap [59].
3.2.2 Estimating NDE and NIE via TMLE
The parametric structural equations we specified simplify the estimation of the NIE and NDE, but they may be restrictive and might not accurately reflect the true data-generating mechanism. TMLE relaxes these linear assumptions and instead identifies the causal effects based on the NPSEM. We give the steps for estimating NDE via TMLE introduced by [68], while the implementation for NIE can be derived as an analog. Without assuming the parametric assumptions, we instead arrive at the following NPSEM
$U=(U_{\boldsymbol{W}},U_{A},U_{M},U_{Y})\sim P_{U}$
$\boldsymbol{W}=f_{\boldsymbol{W}}(U_{\boldsymbol{W}})$
$A=f_{A}(\boldsymbol{W},U_{A})$
$M=f_{M}(\boldsymbol{W},A,U_{M})$
$Y=f_{Y}(\boldsymbol{W},A,M,U_{Y}),$
where $U$ is the unobserved exogenous variables. The likelihood function can be rewritten as
$\mathcal{L}(O)=\mathbb{P}_{\boldsymbol{W}}(\boldsymbol{W})\mathbb{P}_{A}(A|\boldsymbol{W})\mathbb{P}_{M}(M|\boldsymbol{W},A)\mathbb{P}_{Y}(Y|\boldsymbol{W},A,M),$
where we denote $\bar{Q}_{Y}(A,M,\boldsymbol{W})=\mathbb{E}(Y|A,M,\boldsymbol{W})$, $Q_{M}(M|A,\boldsymbol{W})=\mathbb{P}_{M}(M|A,\boldsymbol{W})$, $g(A|\boldsymbol{W})=\mathbb{P}_{A}(A|\boldsymbol{W})$, $q_{\boldsymbol{W}}(\boldsymbol{W})=\mathbb{P}_{\boldsymbol{W}}(\boldsymbol{W})$. TMLE for NDE targets $\bar{Q}_{Y}$ and the mediated mean outcome difference $\mathbb{E}_{Q_{M}}(\bar{Q}_{Y}(1,M,\boldsymbol{W})-\bar{Q}_{Y}(0,M,\boldsymbol{W})\mid\boldsymbol{W},A=0)$, using loss functions and submodels to minimize empirical risk while solving the EIF. We also assume that $\bar{Q}_{Y}$ and $\mathbb{E}_{Q_{M}}$ are bounded in between $[0,1]$. For targeting $\bar{Q}_{Y}$, we minimize the empirical cross-entropy risk for $\bar{Q}_{Y}$:
$\hat{\epsilon}_{1}=\arg\min_{\epsilon_{1}}\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{Y}^{\epsilon_{1}}(O_{i}\mid\hat{Q}_{M},\hat{g}),\quad\mathcal{L}_{Y}^{\epsilon_{1}}(O):=-\left[Y\log\bar{Q}_{Y}^{\epsilon_{1}}(A,M,\boldsymbol{W})+(1-Y)\log\left(1-\bar{Q}_{Y}^{\epsilon_{1}}(A,M,\boldsymbol{W})\right)\right],$
via the logistic working submodel
$\bar{Q}_{Y}^{\epsilon_{1}}(A,M,\boldsymbol{W})=\text{expit}\left\{\text{logit}\left[\hat{\bar{Q}}_{Y}(A,M,\boldsymbol{W})\right]+\epsilon_{1}\,C_{Y}(\hat{Q}_{M},\hat{g};A,M,\boldsymbol{W})\right\},$
where the clever covariate is
$C_{Y}(\hat{Q}_{M},\hat{g};A,M,\boldsymbol{W})=\frac{\mathbf{1}(A=1)}{\hat{g}(1\mid\boldsymbol{W})}\cdot\frac{\hat{Q}_{M}(M\mid 0,\boldsymbol{W})}{\hat{Q}_{M}(M\mid 1,\boldsymbol{W})}-\frac{\mathbf{1}(A=0)}{\hat{g}(0\mid\boldsymbol{W})}.$
The update can be done via $\hat{\bar{Q}}_{Y}^{*}(A,M,\boldsymbol{W})=\bar{Q}_{Y}^{\hat{\epsilon}_{1}}(A,M,\boldsymbol{W}).$ Once we obtain the targeted estimator for $\bar{Q}_{Y}$, we can obtain $\hat{\mathbb{E}}_{M}(\hat{\bar{Q}}_{Y}^{*}|\boldsymbol{W},0)$ as the initial plug-in estimator of the mediated difference in the outcome. The targeting step for $\mathbb{E}_{M}$ uses the proposed loss function
$\mathcal{L}^{\epsilon_{2}}_{M}(\mathbb{E}_{M})=-\mathbf{1}(A=0)[\bar{Q}_{Y}\log\mathbb{E}^{\epsilon_{2}}_{M}+(1-\bar{Q}_{Y})\log(1-\mathbb{E}^{\epsilon_{2}}_{M})].$
Thus, the fluctuation parameter could be computed by
$\hat{\epsilon}_{2}=\arg\min_{\epsilon_{2}}\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}^{\epsilon_{2}}_{M}(\hat{\mathbb{E}}_{M}(\hat{\bar{Q}}_{Y}^{*}|\boldsymbol{W}_{i},0)),$
[…]
$\hat{\psi}^{*}_{\text{NDE}}=\frac{1}{n}\sum_{i=1}^{n}\hat{\mathbb{E}}_{M}^{*}(\boldsymbol{W}_{i}),$
which solves the EIF and attains asymptotic efficiency under regularity conditions, including consistency of the nuisance estimators for $\bar{Q}_{Y}$, $g$, and $Q_{M}$, achieving the semiparametric efficiency bound $\mathrm{Var}(D^{*}(P_{0}))$. For software availability, [47] provides a R package called medoutcon for estimating NIE and NDE using TMLE with the super learners implemented.
4 Simulation Studies
To demonstrate the equivalence of SEM and TMLE under correct model specification and to compare their performance, we conducted simulation studies evaluating both correct and misspecified model scenarios. Both approaches are assessed based on the point estimations and the corresponding confidence set, where TMLE incorporates the Super Learner algorithm.
4.0.1 Methods for Average Treatment Effects
We consider a similar data-generating process in [53], which is equivalent to the following DAG shown in Figure 4.
$W_{1}$$W_{2}$$W_{3}$$W_{4}$AAYY
Figure 5: A structural causal graph adapted from [53] illustrating the relationships between exogenous variables $W_{1}-W_{4}$, binary treatment $A$ and the outcome variable $Y$.
Specifically, the data-generating model is
$W_{1}\sim\text{Bernoulli}(0.5);\quad W_{2}\sim\text{Bernoulli}(0.65)$
$W_{3}\sim\text{Uniform}(\{0,1,2,3,4\});\quad W_{4}\sim\text{Uniform}(\{0,1,2,3,4,5\})$
$A|\boldsymbol{W}\sim\text{Bernoulli}(\text{plogis}(-2.5+0.05\times W2+0.25\times W_{3}+0.6\times W_{4}+0.4\times W_{2}W_{4}))$
$Y|A,\boldsymbol{W}\sim\mathcal{N}\left(-1+\psi\times A+0.1\times W_{1}+0.35\times W_{2}+0.25\times W_{3}+0.2\times W_{4}+3.0\times W_{2}\times W_{4},1\right),$
where $\psi\in\{0.5,1.5\}$ is the true ATE. To showcase the robustness of the TMLE, we consider the following model misspecification:
NoInteraction: The data-generating model remains the same while the interaction terms in the outcome model are missing when calling the functions for computing the ATE.
NonLinear: The data-generating model is modified so that the baseline covariates $\boldsymbol{W}$ and $Y$ are not linearly related, but a high-order polynomial term
$Y|A,\boldsymbol{W}\sim\mathcal{N}\left(-1+\psi A+0.1\cdot W_{1}+0.35\cdot W_{2}+0.25\cdot W_{3}^{4}+0.2\cdot W_{4}^{4}+3.0\cdot W_{2}\cdot W_{4},1\right),$
is added.
NonNormal: The outcome model in the data-generating process is not truly normally distributed, but a student-$t$ with degrees of freedom equal to two. The mean function is still correctly assumed to be linear.
We conducted a Monte Carlo simulation with $n_{\text{sim}}=1000$ with various sample size $n\in\{200,500,750,1000,1500,2000,3000\}$ using both lm and tmle[43] in R and assess the consistency and efficiency using (i) Relative Bias: $n_{\text{sim}}^{-1}\sum_{i=1}^{n_{\text{sim}}}\frac{\hat{\psi}-\psi}{\psi}$; (ii) $95\%$-CI coverage: $n_{\text{sim}}^{-1}\sum_{i=1}^{n}\mathbf{1}_{\psi\in(\ell,u)}$, where $\ell,u$ denote the lower and upper bounds of $95\%$-CI; (iii) Statistical Power against $\psi=0$: $n_{\text{sim}}^{-1}\sum_{i=1}^{n_{\text{sim}}}\mathbf{1}_{0\not\in(\ell,u)}$; (iv) Standardized RMSE: $\sqrt{n_{\text{sim}}^{-1}\sum_{i=1}^{n_{\text{sim}}}\left(\frac{\hat{\psi}-\psi}{\psi}\right)^{2}}$. Instead of linear regression, one could also use the SEM with a logit link to estimate the path coefficient. The superlearner libraries used for tmle are Generalized Linear Model(glm, glm.interaction); generalized additive model(gam); random forest(ranger) and eXtreme Gradient Boosting(xgboost). We then repeat the experiment for CATE, by adding an interaction term between $A$ and $W_{1}$
$Y|A,\boldsymbol{W}\sim\mathcal{N}\left(-1+\psi A+0.5A\times W_{1}+0.1\times W_{1}+0.35\times W_{2}+0.25\times W_{3}+0.2\times W_{4}+3.0\cdot W_{2}\times W_{4},1\right),$
and assess the treatment effects among the group $W_{1}=1$.
4.0.2 Methods for Mediation Analysis
To conduct the simulation studies for mediation analysis, we consider the following data-generating process
$W\sim\mathcal{N}(0,1)$
$A|W\sim\text{Bernoulli}(\text{plogis}(0.5\times W))$
$M|A,W\sim\mathcal{N}(A+0.5\times W,1)$
$Y|A,W,M\sim\mathcal{N}(2\times A+M+0.8\times W,1).$
The mediation effects under this data-generating process could be identified to be $\psi_{\text{NIE}}=1$ and $\psi_{\text{NDE}}=2$. To conduct the mediation analysis, both SEM and TMLE are fitted and assessed. SEM was implemented using lavaan in R, and the confidence interval with the corresponding inference were based on bootstrap with a number of samples equal to 1000. The TMLE for the mediation analysis was implemented using medoutcon [47] with Generalized Linear Models and Random Forest as the superlearner algorithms being used to obtain the estimates in each of the steps. Due to the computational costs, we conduct a Monte Carlo simulation with $n_{\text{sim}}=200$ with sample sizes $n\in\{500,800,1000,1500,2000,2500,3000,5000\}$ and assess both of the estimators using the same metrics that were computed for ATE and CATE.
For the model misspecification, we consider the following two scenarios: MisspecYW: the outcome path is modified to be non-linear such that
$Y|A,W,M\sim\mathcal{N}(2\times A+M+0.8\times W^{4},1).$
MisspecMWYW: both the mediation path and outcome path are misspecified such that
$M|A,W\sim\mathcal{N}(A+0.5\times W^{2},1)$
$Y|A,W,M\sim\mathcal{N}(2\times A+M+0.8\times W^{4},1).$
and then re-fit both approaches to the misspecified cases to compare the performance between SEM and TMLE.
4.0.3 Results for Average Treatment Effects
Figures 6 and Figure 8 show the distributional boxplot across $1000$ Monte Carlo simulations for ATE and CATE, respectively. When the outcome models are correctly specified, both the Linear Regression and TMLE result in estimates that are centred around the true values. However, if the functional form of the outcome models is misspecified, Linear Regression failed to result in an estimation centred at the true values even under a large sample size $n=3000$, while TMLE managed to achieve this even when the sample size is relatively small $n=500$. This is expected as the OLS is no longer unbiased when the model is misspecified, but TMLE is robust to model misspecification by the targeting steps and the incorporation with the super learner. This pattern was consistently observed for both ATE and CATE.
Figure 6: Distributional Boxplot for ATE across 1000 Monte Carlo simulations.
Figure 7: Performance of Regression and TMLE for ATE under various metrics against different sample sizes.
Figure 7 and Figure 9 show the performance of linear regression and TMLE under different metrics for estimating ATE and CATE, respectively. As expected, we observe that the relative bias and RMSE of TMLE and Linear Regression are close to zero when the model is correctly specified, indicating unbiasedness.
Figure 8: Distributional Boxplot for CATE across 1000 Monte Carlo simulations.
Figure 9: Performance of Regression and TMLE for CATE under various metrics against different sample sizes.
When models are misspecified, the regression approach yields biased estimates even with a relatively large sample size of $n=5000$, whereas TMLE produces unbiased estimates. This is expected as the regression converges to the pseudo-true parameter. Under nonlinear relationships, the bias for TMLE approaches zero at $n=500$ and remains unbiased thereafter. The statistical power against $H_{0}:\psi=0$ converges to one for both linear regression and TMLE when the model is correctly specified. However, when the outcome model is not truly linear, linear regression exhibits power significantly below one due to its biased estimation, while TMLE’s power approaches one at around $n=500$ and remains high thereafter. The coverage of 95% confidence intervals follows a similar pattern: linear regression shows a downward trend approaching zero under misspecification, whereas TMLE maintains coverage close to $0.95$. Moreover, although TMLE relies on asymptotic properties, it achieves the desired CI coverage and statistical power relatively quickly as the sample size increases. When the normality assumption is violated, TMLE performed slightly better than the regression, but the difference is not significant. These patterns hold consistently for both ATE and CATE, demonstrating TMLE’s robustness to model misspecification.
4.1 Results for Mediation Analysis
Figure 10 presents the performance of SEM and Targeted TMLE across key statistical metrics, including bias, root mean square error (RMSE), $0.95$-confidence sets coverage probability, and power against $\psi=0$, as influenced by sample size and model specification. When all structural pathways are correctly specified as linear, both SEM and TMLE produce unbiased estimates with RMSE nearing zero, demonstrating their equivalence in the absence of model misspecification. This aligns with their theoretical properties when assumptions hold.
However, significant differences emerge under both model misspecifications. SEM generates biased estimates for the NDE when the YWY-W path is no longer linear, and its confidence intervals fail to achieve the expected $0.95$ coverage, indicating a loss of validity due to its dependence on correct functional forms. The NIE is also affected since SEM maximizes the joint likelihood induced by the model. In contrast, TMLE remains robust, delivering unbiased NDE estimates and maintaining valid confidence intervals even with a sample size as small as $n=500$.
The disparity becomes more pronounced when both the YWY-W and MWM-W (mediator-exposure) pathways are misspecified. SEM yields biased estimates for both NDE and NIE, with confidence intervals that no longer provide the nominal $0.95$ coverage, likely due to its vulnerability to multiple model misspecifications. TMLE, however, continues to offer unbiased estimates for both effects, with desired confidence sets that maintain $0.95$ coverage. This robustness is expected for TMLE, owing to its double robustness properties and integration with machine learning algorithms via the Super Learner.
Figure 10: Performance of SEM and TMLE for NDE, NIE, and TE under various metrics against different sample sizes.
In terms of the power analysis, TMLE consistently surpasses SEM in power under misspecified models, indicating greater efficiency in detecting true effects. This suggests that TMLE is statistically more robust than parametric SEM for mediation analysis, especially when the data-generating mechanism is uncertain or incompletely specified.
5 Applications
To demonstrate the practical utility of TMLE in estimating ATE, CATE, and mediational effects, we apply it to a real-world sociological question: the causal impact of multidimensional poverty on rural children’s access to high school education in China, with a focus on gender heterogeneity and the mediating role of educational expenditure.
5.1 Data and Study Design
The analysis draws from six waves between 2012 and 2022 of the China Family Panel Study (CFPS), a nationally representative longitudinal survey. The sample consists of $2881$ rural children aged from six to fifteen in the 2012 wave and newly sampled children aged from eight to fifteen in the 2014 wave, tracked for high school enrollment at ages 15 or 16 in subsequent waves. The binary outcome is high school enrollment[50]. Following [62], multidimensional poverty was measured using nine indicators across three dimensions: health (nutrition, child mortality), education (parental years of schooling, children’s elementary school attendance), and living standards (cooking fuel, sanitation, drinking water, electricity, housing assets). Each dimension and its indicators within each dimension were equally weighted. A child was considered multidimensionally poor if their weighted deprivation score was 0.33 or higher. To satisfy the unconfoundedness assumption, covariates include individual factors (age, gender, siblings), family characteristics (household registration, living arrangements, parental education and occupation, political identity, $\log$-per capita income, Dibao receipt), school quality, and province fixed effects. All the above covariates were drawn from the first wave for each sample. Additionally, we selected children’s educational expenditure as a mediator, measured in two ways: first, as the logarithm of the absolute expenditure amount; second, as the proportion of total household income spent on education.
5.2 Estimation of (Conditional)Average Treatment Effects
To assess the average impact of multidimensional poverty on children’s access to high school, parametric logistic regression using glm in R, and TMLE with default superlearners in tmle, including Regularized GLM and Bayesian Additive Regression Trees [43] were implemented.
Table 1: The Average Impact of Multidimensional Poverty on Children’s Access to High School.
Model 1 Model 2
Method TMLE LOGIT
     ATE -0.106 -0.140
     S.E. 0.030 0.046
     95% Confidence Interval (-0.164, -0.048) (-0.230, -0.051)
The overall impact of multidimensional poverty on high school access is presented in Table 1. The TMLE estimated an ATE of -0.106 (95% CI: -0.164 to -0.048). The logistic regression yielded a larger ATE of -0.140 (95% CI: -0.230 to -0.051). Both approaches reported negative effects of multidimensional poverty on high school access, with statistical significance at $\alpha=0.05$.
Table 2: The Average Impact of Multidimensional Poverty on Children’s Access to High School by Gender.
Model 3 Model 4 Model 5 Model 6
Method TMLE TMLE LOGIT LOGIT
     Subpopulation Male Female Male Female
     ATE -0.169 -0.084 -0.147 -0.134
     S.E. 0.041 0.035 0.066 0.063
     95% Confidence Interval (-0.249, -0.089) (-0.153, -0.016) (-0.275, -0.018) (-0.256, -0.011)
Table 2 reports the heterogeneous effects by gender, where the ATE is separately estimated for male and female subpopulations. For males, the TMLE estimated an ATE of -0.169 (95% CI: -0.249 to -0.089), while the logistic regression estimated an ATE of -0.147 (95% CI: -0.275 to -0.018). For females, the TMLE indicated a smaller ATE of -0.084 (95% CI: -0.153 to -0.016), and the logistic regression estimated an ATE of -0.134 (95% CI: -0.256 to -0.011). Both methods consistently indicate a larger negative impact on males compared to females, with TMLE estimates showing a more pronounced gender difference. The effects are statistically significant under $\alpha=0.05$ for both female and male under both approaches.
This aligns with the previous research, which showed in the education stage, boys are often more sensitive to familial disadvantages [35, 54, 51]. Families often impose higher educational expectations on boys, reflecting broader societal norms that link male identity to achievement. In the context of poverty, however, limited economic resources and constrained parental support can impede boys’ capacity to meet these expectations, exacerbating vulnerabilities and leading to lower academic attainment, diminished motivation, and increased risk of adverse educational outcomes relative to girls.
5.3 Mediation Analysis
Beyond estimating (conditional) average treatment effects, sociologists are often equally concerned with the mechanisms through which causal relationships unfold. To assess the direct and indirect effects, SEM using lavaan [60] and TMLE using medoutcon [47] in R were implemented. To account for the potential non-linearity and sparsity, algorithms including Random Forest, LASSO, Gradient Boosting, and Generalized Additive Model are specified in the superlearner when implementing the TMLE.
Table 3: Mediating Effect Decomposition of Multidimensional Poverty on Children’s Access to High School.
Indirect Effect Direct Effect
Standardized Effect 95%-CI Standardized Effect 95%-CI
Model 7(SEM) -0.023 (-0.037, -0.009) -0.097 (-0.187, -0.007)
Model 8(SEM) -0.020 (-0.034, -0.006) -0.100 (-0.192, -0.008)
Model 9(TMLE) -0.110 (-0.216, -0.003) -0.150 (-0.302, 0.002)
Model 10(TMLE) -0.105 (-0.198, -0.013) -0.142 (-0.296, 0.011)
Table 3 presents the results of the mediation analysis. For SEM with educational expenditure measured as the logarithm of absolute spending, the direct effect is -0.097 (95%-CI: -0.187 to -0.007) and the indirect effect is -0.023 (95%-CI: -0.037 to -0.009). For SEM using the ratio of educational expenditure to household income, the direct effect is -0.100 (95%-CI: -0.192 to -0.008) and the indirect effect is -0.020 (95%-CI: -0.034 to -0.006). For TMLE with educational expenditure measured as the logarithm of absolute spending, the direct effect is -0.150 (95%-CI: -0.302 to 0.002) and the indirect effect is -0.110 (95%-CI: -0.216 to -0.003). For TMLE using the ratio of educational expenditure to household income, the direct effect is -0.142 (95%-CI: -0.296 to 0.011) and the indirect effect is -0.105 (95%-CI: -0.198 to -0.013). Though the indirect effects obtained by TMLE remain statistically significant, the direct effects are no longer significant. This contradicts SEM, potentially due to model misspecification.
6 Discussion
This paper clarifies the conceptual and practical links between Structural Equation Modeling (SEM) and Targeted Maximum Likelihood Estimation (TMLE) for causal inference. We framed path coefficients in SEM and identified functionals in nonparametric structural equation models (NPSEM) as two views of the same goal: estimating causal parameters (ATE, CATE, direct/indirect effects) under explicit causal assumptions. Despite their shared objectives, we compare the two approaches across diverse settings, including correct model specification and misspecification. Finally, we demonstrate their practical utility using real-world data to assess the causal effects of poverty on high school access.
6.1 Key findings
Our simulation studies and empirical research yield three primary findings. First, when the outcome and treatment models are correctly specified and identification assumptions hold, linear SEM and TMLE deliver comparable point estimates and valid uncertainty quantification for ATE, CATE, and mediation. Second, under misspecification—omitted interactions, nonlinearities, or non-normal outcomes—TMLE retained consistency and near-nominal coverage through double robustness and targeting via the efficient influence function, while SEM-based linear estimators exhibited bias and degraded coverage. Third, in the mediation setting, TMLE maintained valid intervals when the outcome mechanism departed from linearity; SEM, in turn was sensitive to such deviations. These results position TMLE as a pragmatic, misspecification-resilient complement to SEM.
6.2 Scope and limitations
Our SEM discussion is intentionally limited to path analysis with observed variables. We do not treat latent-variable SEM (i.e., measurement models, factor loadings, and their implications for identification, reliability, and attenuation). As a result, our comparisons do not cover (i) bias introduced by measurement error in indicators of $A$, $M$, or $Y$; (ii) the interplay between measurement model misspecification and structural paths; or (iii) methods that integrate targeted learning with latent-variable measurement (e.g., SEM with latent variables or any TMLE extensions for including latent variables). Extending the analysis to latent constructs is a valuable direction for future work, especially in domains where psychological or sociological constructs are measured with multi-item scales.
6.3 When to prefer which approach
SEM remains attractive when (i) measurement models for latent constructs are central; (ii) theory supports low-dimensional linear structures; and (iii) transparent parametric constraints are part of the scientific question (e.g., equality constraints, cross-loading tests). TMLE is preferable when (i) the functional form is uncertain, (ii) effect heterogeneity (CATE) is expected, (iii) mediators and outcomes exhibit nonlinearities or interactions, or (iv) valid coverage is prioritized under plausible misspecification.
6.4 Extensions and open directions
Several extensions merit attention. First, longitudinal treatments and mediators call for sequential TMLE with time-varying confounding and stochastic interventions, whereas SEM analogs rely on cross-lagged or latent growth structures; a formal comparison is warranted. Second, missing data can be integrated with targeted learning via inverse-probability weighting and augmented estimators; SEM users may consider full-information ML but should assess MNAR risks. Third, interference and spillovers (common in social settings) violate SUTVA; recent network-TMLE developments could be adapted. Fourth, transportability across populations suggests integrating TMLE with reweighting for different covariate supports; SEM analogs could impose equality constraints across groups while allowing distributional shifts. Finally, design-based targeted learning (e.g., targeted regularization, collaborative TMLE, debiased machine learning) may further stabilize small-sample performance.
6.5 Conclusion
SEM and TMLE are complementary tools for causal inference. When theory-driven linear structures are credible, SEM remains efficient and interpretable. When misspecification is a live concern—as it often is in applied work—TMLE offers robustness and valid inference by targeting the estimand directly. Placing identification upfront, followed by either a well-specified SEM (including, in future work, measurement models for latent constructs) or a targeted learning pipeline with flexible learners, can substantially improve the credibility of causal conclusions in the social and health sciences.